OpenStudy (anonymous):

one-step inequalities. Think of a situation in your day where you find a similar situation. Complete the five steps below •Describe the every-day situation you are solving. •Create a verbal model. •Define a variable. •Substitute the known amounts and variable. •Solve the inequality.

OpenStudy (anonymous):

your trying to save up some money for a present for your mom. The present can't be more then $200. You already have $80 saved and you get $20 a week. How many weeks will it take you to save up the rest of the money ? verbal model : 80 + 20w <= 200 define the variable : the variable is w and it represents the number of weeks solve : 80 + 20w <= 200 20w <= 200 - 80 20w <= 120 w <= 120/20 w <= 6